# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ 2.6 Performance Evaluation
The variety of datasets requires selecting suitable evaluation metrics to fairly compare and evaluate prediction methods. As discussed in the Section 2.5, unimodal methods predict a single deterministic trajectory or the parameters of a unimodal trajectory distribution. In contrast, multimodal methods generate multiple possible trajectories or a multimodal probability distribution. Below, we review evaluation metrics for both unimodal and multimodal methods, where $\zeta_{i,t}$ and $\hat{\zeta}_{i,t}$ denote the predicted and ground truth trajectories for the i-th agent at time step $t$, respectively.
@@
  • 5.
    Miss Rate (MR) measures the ratio of cases where the predicted trajectory deviates significantly from the ground truth, based on criteria such as a final displacement threshold (e.g., greater than 2 meters).
Lastly, measuring and reporting inference time is crucial for a thorough evaluation of prediction algorithms, as it reflects their suitability for real-time operation. To compute the reported metrics accurately, alignment between predicted and ground truth trajectories is required. In the absence of explicit object identifiers, this alignment can be performed using proximity-based matching techniques, such as IoU thresholds or nearest centroid matching. Once the alignment is completed, the reported metrics can be calculated to assess prediction accuracy.
@@ 2.7 Datasets and Leaderboard
Assessing the effectiveness of trajectory prediction models involves not only selecting accuracy metrics but also designing diverse test scenarios to evaluate the models’ performance under various factors, such as the number of interacting agents and the degree of maneuvering in the data. The test scenarios can vary widely in complexity, location, and traffic density, including intersections, highways, mixed urban streets, crowded public spaces, residential areas, and more. Each scenario presents unique challenges and variability that the algorithm must handle effectively to demonstrate its practicality and potential for real-world autonomous driving applications. Table 2 lists trajectory prediction datasets that can be utilized to test prediction models. It provides details on the modalities used to collect the datasets, along with short descriptions and leaderboard solutions. For some datasets, certain details are omitted, encouraging readers to refer to the datasets’ official websites for the most up-to-date information.
@@ 2.8 Prediction Paradigms
Figure 6: Overview of prediction paradigms in autonomous driving: (a) Detect-Track-Predict: a modular pipeline where detections are used to initialize and update active tracklets, which are subsequently provided as past trajectories to the prediction module; (b) End-to-End Detection-Prediction: an integrated approach where detections are pooled and directly utilized by the prediction module; (c) Query-Based Prediction: a query-driven framework where predictions are refined based on object queries applied to encoded sensor data.
This section discusses three existing prediction paradigms: Detect-Track-Predict, End-to-End Approaches, and Query-Based Approaches, which are graphically illustrated in Figure 6. For each paradigm, we outline the general concept and highlight its advantages and limitations for integration.
@@ 2.8.1 Detect-Track-Predict
The Detect-Track-Predict paradigm is a modular cascade approach consisting of detection, tracking, and prediction modules. First, the detection module localizes objects from sensor data. Next, the tracking module initializes tracklets for the detected objects and maintains their trajectories consistently over time. Finally, the prediction module takes the active tracklets and forecasts their future positions. This approach is one of the most common implementations in autonomous driving and is broadly used in the literature. Its modular design makes it easier to debug, as most datasets provide ground truth annotations for each of the modules. Additionally, the extensive research output for detectors [33] and trackers [188], covering both 2D and 3D modalities, further supports its seamless development.
The latest detection approaches leverage data fusion from cameras and LiDAR sensors to enhance detection and tracking accuracy as shown in Figure 7. Early Fusion integrates raw sensor data, such as images and point clouds. A common implementation involves detecting objects in the image, projecting the point cloud onto the 2D image plane, and matching the projected points with the detected objects to classify them and measure distances accurately. Guidolini et al. [189] proposed such a pipeline for combining the strengths of both modalities. In contrast, Late Fusion operates at the detection stage, combining results from LiDAR and camera detections. This can be implemented in two ways: (1) merging 3D bounding boxes detected independently from LiDAR and camera, or (2) projecting LiDAR’s 3D bounding boxes onto the 2D image plane and aligning them with the 2D detections from the camera. This strategy provides flexibility, allowing fusion to occur in either the two-dimensional or three-dimensional space, depending on the accuracy of the detection techniques for both modalities. Another fusion strategy, Augmenting Point Clouds, enhances LiDAR data by projecting point clouds onto images and enriching each point with matching image features. This augmented representation is then fed into a LiDAR-only detection method. Vora et al. [190] proposed PointPainting, which projects LiDAR points onto semantic segmentation outputs from the camera and appends the corresponding class scores to each point. Camera-to-LiDAR projections can result in a loss of rich semantic information. Bird’s-eye view (BEV) Fusion mitigates this issue by combining features from different modalities within a unified BEV representation, providing a more comprehensive spatial understanding, but resulting in a prolonged inference time. For a detailed overview, we refer the reader to the multi-sensor fusion framework presented in [191].
@@
Kalman Filter-based trackers [192, 193, 194], leverage predicted object states to mitigate detection inaccuracies caused by sensor noise or missed detections. Advanced trackers [195] address partial occlusions, ensuring robust trajectory maintenance and minimizing false positives that lead to ghost tracklets. Prediction models rely on the precision of trajectories over the past 10–20 timestamps, making robust tracklets essential for accurate forecasting. Excessive ghost tracklets result in non-existent predictions being passed to the planner, potentially degrading overall system performance.
@@
2.8.2 End-to-End Models
End-to-End models [167, 196, 197, 198, 100] are purposed to address the limitations of modular approaches by jointly optimizing object detection and trajectory prediction within a single model. This approach assume no explicit tracker and capturing temporal dependencies directly from raw sensor inputs, e.g., sequential LiDAR or camera frames. As shown in Figure 6, a pool module can be integrated to bridge the gap between detection and prediction. The pool aggregates and encodes object features from detection outputs into a unified representation optimized for trajectory prediction. Additionally, it organizes and streamlines inputs to the predictor, dynamically filtering or prioritizing detections based on their relevance to reduce redundant or irrelevant information.
By sharing high-dimensional scene features, end-to-end models retain richer contextual information, enabling the learning of intermediate representations that benefit both detection and prediction tasks. This paradigm avoids the information bottlenecks of cascade approaches by eliminating the tracking module and jointly optimizing detection and prediction. However, despite these advantages, errors in the shared features from the detection stage can propagate through the system, affecting overall performance.
@@ 2.8.3 Query-Based Prediction Models
Query-Based prediction models represent a cutting-edge approach to trajectory prediction by leveraging learnable queries within a unified framework. Rather than splitting detection and prediction into separate stages, query-based models directly focus on objects of interest, iteratively refining their trajectory predictions. These learnable object queries encapsulate spatial and temporal parameters, guiding the model in producing consistent predictions. By integrating contextual data over multiple refinement steps, query-based models capture fine-grained details and interaction dynamics. A key advantage of this approach is the elimination of error propagation across separate subsystems, as predictions are generated holistically within a single framework. In the following discussion, we highlight several query-based methods proposed in the literature.
Zhou et al. [179] proposed a query-centric paradigm for trajectory prediction, encoding scene information to reuse encodings efficiently. The model employs anchor-free and anchor-based queries to iteratively refine trajectory proposals, enabling parallel multi-agent trajectory decoding. Next, Hu et al. [199] introduced UniAD, a unified framework integrating perception, prediction, and planning with a query-based design. Modules such as TrackFormer, MapFormer, MotionFormer, and OccFormer perform tasks like detection, tracking, road segmentation, future movement prediction, and occupancy forecasting. The final planner predicts waypoints while avoiding collisions and optimizing trajectories, ensuring coordinated decision-making. Its scene-centric approach considers agent interactions and environmental context. Casas et al. [200] presented DeTra, a query-based framework unifying detection and trajectory prediction. Temporal and spatial object queries are used to process LiDAR point clouds and HD maps, iteratively refining object poses. Lastly, Agro et al. [166] proposed a query-based occupancy flow prediction model where query points represent BEV spatio-temporal coordinates. These guide the network to predict occupancy probabilities and motion flow dynamically, integrating past and present spatial information.
@@ 3 Survey of Modelling Methods
This section describes the refined taxonomy of trajectory prediction approaches. To systematically categorize the wide range of prediction models in the literature, the taxonomy is structured based on modeling methods. It builds upon the categorizations presented in reviews [4, 5], and is further extended based on a review of recent works in which prediction models were systematically tagged and grouped.
@@ 4 Incorporating Driving Knowledge
To enhance performance and increase the explainability of trajectory prediction frameworks, recent studies [476, 477, 478, 479] have incorporated driving knowledge. The underlying intuition is that integrating external information to constrain motion reduces the set of possible future trajectories. In most scenarios, vehicles operate within designated lanes and adhere to traffic rules. Including information such as lane geometry, routes, road boundaries, and traffic regulations at intersections and roundabouts plays a crucial role in accurate predictions. Another type of driving knowledge that can improve prediction accuracy includes vehicle motion constraints, such as turning radius and braking distance, as well as pedestrian mobility limitations. Additionally, accounting for different driving styles and behavioral priors can improve model generalization and performance. For instance, in some regions, it is common to turn without signaling or to take a turn from a neighboring lane instead of the designated one to avoid traffic congestion. While such actions may constitute traffic rule violations, they can also be viewed as region-specific behavior patterns, reflecting common trends rather than individual choices. This section explores different types of driving knowledge and the ways in which this knowledge is incorporated into prediction frameworks. For a more detailed discussion, the reader is encouraged to refer to the survey paper [10].
@@ 4.1 Types of Driving Knowledge
We categorize driving knowledge into three groups: i) map-related information encoding the static environment, such as road topology, pedestrian crossings, and stop lines; ii) agent-related information about the kinematics of vehicles and the physical constraints of pedestrians; and iii) traffic-related knowledge, which indicate any additional information related to traffic rules, such as driving styles and interaction patterns between road users.
@@ 4.1.1 Map-related knowledge
HD maps provide information about road geometry, traffic lights, stop lines, pedestrian crosswalks, and traffic signs. Section 2.2 discusses rasterizing map data into grid images and encoding them using CNNs. While effective for extracting spatial features, this method often results in information loss, as the structured and geometric details of the map are rasterized, reducing precision. Recent advancements use GNNs to encode HD maps, preserving structural and semantic details while remaining scalable and effective.
One spotlight example is VectorNet, proposed by Gao et al. [407], which represents HD map elements and agent trajectories in a vectorized format. Each map element is broken into polylines composed of individual vectors, forming local graphs where nodes represent vectors and edges capture spatial or sequential relationships. These local graphs are aggregated into a global graph to encode lane-to-lane or agent-to-lane interactions using a hierarchical GNN. This approach enables interaction-aware feature encoding for downstream tasks. Next, Liang et al. [374] introduced LaneGCN, which leverages HD maps by constructing a lane graph where each lane segment is a node, and edges denote connections like predecessors, successors, and neighboring lanes. The model dynamically encodes the scene by capturing the topology and dependencies of the lane graph. By integrating actor-map interactions through a fusion network, it combines real-time traffic information with map-based context for prediction. Deo et al. [480] proposed a complementary approach, which focuses on conditioning trajectory predictions on specific paths within the lane graph. The proposed method uses discrete policy rollouts to select plausible routes, reducing computational complexity by targeting relevant map regions. By decoupling lateral variability (e.g., route choices) and longitudinal variability (e.g., speed, acceleration). Focusing predictions on sampled paths, the model enhances multimodal diversity, scene compliance, and trajectory prediction.
@@ 4.1.2 Agent-related knowledge
Agent-related knowledge enables prediction models to account for behavior and motion constraints. A vehicle’s kinematics determine its turning radius, braking distance, and acceleration capacity — all factors that significantly impact its potential future trajectory [218, 481]. Another aspect of agent-related driving knowledge involves understanding different driving styles. By analyzing a target vehicle’s historical trajectories, models can draw on established principles of driving behavior to identify patterns that align with known driving styles. Observing maneuvers and typical trajectory choices enables the model to estimate the driver’s style based on these predefined categories. For instance, an aggressive driving style might suggest a higher likelihood of selecting risky trajectories, such as overtaking with insufficient gaps or increasing the risk of collisions by forcing following vehicles to brake abruptly.
@@ 4.1.3 Traffic-related knowledge
Traffic-related knowledge encompasses additional driving information beyond agent- and map-related categories. Traffic rules provide a foundation for constraining vehicle behaviour and narrowing the set of plausible trajectories. Incorporating explicit traffic rules, such as speed limits and right-of-way regulations, into prediction models helps ensure safety and compliance with real-world driving norms. Kumar et al. [482] discuss the impact of traffic rules on intersection interactions, where vehicles may be required to stop, yield, or proceed depending on traffic signals or signs. In roundabouts, Hu et al. [483] highlight how right-of-way rules help define permissible behaviors and generate conflict-free trajectories for all participants. Similarly, Ju et al. [484] note that vehicle acceleration often reflects adherence to traffic regulations, such as slowing down at crosswalks or maintaining appropriate speeds in speed-restricted zones. By encoding traffic rules into models, predictions can be better aligned with expected behaviors in regulated environments.
Another type of traffic-related knowledge is interaction knowledge, which helps participants coexist in dynamic environments. In the study [482], pairwise interactions at intersections are categorized into discrete types such as ignoring, going, or yielding, based on the behavior of the target vehicle. Ban et al. [485] extend this approach by introducing follow and yield interactions, requiring models to explicitly predict these labels to better capture traffic dynamics. For more complex scenarios, such as roundabouts, Hu et al. [483] define interaction types using reference paths of interacting vehicles, enabling the generation of conflict-free trajectories. Ju et al. [484] link interactions to acceleration patterns, proposing a sub-network that predicts vehicle acceleration to infer interaction-driven trajectories. Additionally, social norms and informal yielding practices can also be of use, e.g., a driver signaling a pedestrian to pass. These implicit forms of interaction knowledge can be formalized and incorporated into models to better predict complex and multimodal interactions in dense traffic scenarios.

### v3

@@ 2.6 Performance Evaluation
The most common evaluation setting in existing literature for trajectory prediction models is the open-loop mode, where predicted trajectories are evaluated by directly comparing them to ground-truth future paths, without quantifying the impact on downstream modules. Further we review evaluation metrics for both unimodal and multimodal methods discussed in the Section 2.5, where $\zeta_{i,t}$ and $\hat{\zeta}_{i,t}$ denote the predicted and ground truth trajectories for the i-th agent at time step $t$, respectively.
@@
When object IDs are preserved, such as in evaluations using ground-truth tracklets (past trajectories), computing prediction metrics is straightforward. However, in scenarios where ground-truth identifiers are unavailable, predicted and ground-truth objects must first be matched, typically using IoU thresholds or nearest-centroid distance. Only then can accuracy metrics be calculated. In such settings, it is essential to report the Miss Rate (MR), which quantifies the proportion of agents entirely missed by the perception-prediction pipeline and thus never reported to the planner for collision avoidance. Another critical metric that should become standard in benchmarking is inference time, as it directly reflects the model’s suitability for real-time deployment.
In closed-loop evaluation, prediction models are assessed within the full autonomy stack [179], capturing how prediction errors propagate to downstream planning and impact ego-vehicle safety. In this setting, the models can be evaluated from the point of inter-frame prediction consistency, by measuring how frequently the model triggers unnecessary rerouting. Another important factor is the model’s level of conservatism, which determines which predictions should be passed to the planner. Overly conservative models may output fully occupied grids, causing the ego-vehicle to stop unnecessarily, even when feasible paths exist. More on discussion on open and closed-loop evaluations we refer the reader to the study [152].
Finally, for proper cross-dataset evaluation and deployment, it is essential to account for differences in sampling rates to ensure fair performance comparisons. For example, if a model is trained on data sampled at 20 frames per second (FPS) and deployed in a 10 FPS setting, downsampling must be applied. Similarly, if the model is trained to predict trajectories at 5 steps per second, this defines the motion dynamics it learns. To deploy such a model in a 10 Hz setup, resampling or appropriate temporal alignment is required. Some works to avoid these challenges caused by discretizations shift to continuous domains by representing trajectories as continuous stochastic process [180].
@@ 2.7 Datasets and Leaderboard
Assessing the effectiveness of trajectory prediction models require diverse test scenarios to evaluate the models’ performance under various factors. The test scenarios can vary widely in complexity, location, and traffic density, including intersections, highways, mixed urban streets, crowded public spaces, residential areas, and more. Each scenario presents unique challenges and variability that the model must handle effectively to demonstrate its practicality and potential for real-world autonomous driving applications. Table 2 lists trajectory prediction datasets that can be utilized to test forecasting models. It provides details on the modalities used to collect the datasets, along with short descriptions and leaderboard solutions. For some datasets, certain details are omitted, encouraging readers to refer to the datasets’ official websites for the most up-to-date information.
@@ 2.8 Prediction Paradigms
Figure 6: Overview of prediction paradigms in autonomous driving: (a) Detect-Track-Predict: a modular pipeline where detections initialize and update tracklets, which are then passed as past trajectories to the prediction module; (b) Joint Learning Detection–Prediction: an integrated setup where detections are pooled into object-level representations and directly consumed by the prediction module; (c) Query-Based Prediction: a query-driven framework where spatial and temporal object queries attend to encoded scene and map features to jointly produce detections and predictions.
This section discusses three existing prediction paradigms: Detect-Track-Predict, Joint Learning Approaches, and Query-Based Approaches, which are graphically illustrated in Figure 6. For each paradigm, we outline the general concept and highlight its advantages and limitations for integration.
@@ 2.8.1 Detect-Track-Predict
The Detect-Track-Predict paradigm is a modular cascade approach consisting of detection, tracking, and prediction modules. First, the detection module localizes objects from sensor data. Next, the tracking module initializes tracklets for the detected objects and maintains their trajectories consistently over time. Finally, the prediction module takes the active tracklets and forecasts their future positions. This approach is one of the most common implementations in autonomous driving and is broadly used in the literature. Its modular design makes it easier to debug, as most datasets provide ground truth annotations for each of the modules. Additionally, the extensive research output for detectors [35] and trackers [200], covering both 2D and 3D modalities, further supports this paradigm development.
The latest detection approaches leverage data fusion from cameras and LiDAR sensors to enhance detection and tracking accuracy as shown in Figure 7. For early fusion, a common implementation involves detecting objects in the image, projecting the point cloud onto the 2D image plane, and matching the projected points with the detected objects for depth estimation Guidolini et al. [201]. Late Fusion operates at the detection stage, combining LiDAR and camera detection outputs, by projecting LiDAR’s 3D bounding boxes onto the 2D image plane and aligning 2D representation of LiDAR detections with the 2D detections from the camera. Another fusion strategy, Augmenting Point Clouds, enhances LiDAR data by projecting point clouds onto images and enriching each point with matching image features. This augmented representation is then fed into a LiDAR-only detection method. Vora et al. [202] proposed PointPainting, which projects LiDAR points onto semantic segmentation outputs from the camera and appends the corresponding class scores to each point. Camera-to-LiDAR projections can result in a loss of rich semantic information. Bird’s-eye view (BEV) Fusion mitigates this issue by combining features from different modalities within a unified BEV representation, providing a more comprehensive spatial understanding, but resulting in a prolonged inference time, e.g., as detection inference speed of 190 ms for BEVFusion [203]. For a detailed overview, we refer the reader to the multi-sensor fusion framework presented in [204].
@@
The trackers are responsible of maintaining consistent object identities and motion estimates across frames. Kalman Filter-based trackers [205, 206, 207], leverage predicted object states to mitigate detection inaccuracies caused by sensor noise or missed detections. Advanced trackers [208] address partial occlusions, ensuring robust trajectory maintenance and minimizing false positives that lead to ghost tracklets. Prediction models rely on the precision of trajectories over the past 10–20 timestamps, making robust tracklets essential for accurate forecasting. Excessive ghost tracklets result in non-existent predictions being passed to the planner, potentially degrading overall system performance.
@@
2.8.2 Joint Learning Models
Joint learning (perception & prediction) models [209, 42, 210] are designed to optimize object detection and trajectory prediction under a unified training objective. These models can operate without tracker, capturing temporal dependencies directly from sequential sensor inputs frames. As shown in Figure 6, a pool module can be integrated to bridge the gap between detection and prediction. The pool aggregates and encodes object features from detection outputs into a unified representation tailored for trajectory forecasting. It also organizes and streamlines the inputs to the prediction module, dynamically filtering or prioritizing detections based on their relevance to the forecasting task. Training is typically guided by a compound loss over both detection and prediction objectives.
Depending on the design, joint learning can involve different degrees of integration. Some models share intermediate scene features between perception and prediction, enabling contextual reuse across tasks, e.g, FIERY [209] reuses BEV-encoded features for both detection and forecasting. PnPNet [210] supports shared gradient flow across modules, allowing the prediction loss to propagate through the detector and encoder. Alternatively, structurally decoupled modules may be co-optimized via a combined loss, as in FaF [42], which performs detection and forecasting using a shared LiDAR backbone but without full gradient sharing between modules. The limitation of these models is that the tight coupling of modules in joint learning blurs task boundaries, making it challenging to debug and attribute responsibility for poor performance.
@@ 2.8.3 Query-Based Prediction Models
Query-based prediction models [211] unify detection and forecasting by leveraging learnable object queries that are iteratively refined to predict agent trajectories. These queries encode spatial and temporal information and directly attend to scene and map features, removing the need for separate detection and prediction stages. As shown in Figure 6, predictions are generated holistically, reducing error propagation across subsystems.
Zhou et al. [191] propose DenseTNT, which uses anchor-based and anchor-free trajectory queries to refine multi-agent predictions from shared scene encodings. Hu et al. [212] present UniAD, a full-stack framework that uses task-specific queries (e.g., TrackFormer, MotionFormer) to handle detection, tracking, forecasting, and planning jointly. Casas et al. [213] introduce DeTra, where spatial and temporal queries jointly estimate object poses and future motion directly from LiDAR and HD maps. Agro et al. [176] propose a scene-centric model where spatio-temporal BEV queries guide occupancy and flow prediction over time.
@@ 3 Survey of Modelling Methods
This section describes the refined taxonomy of trajectory prediction approaches. To systematically categorize the wide range of prediction models in the literature, the taxonomy is structured based on modeling methods. It builds upon the categorizations presented in reviews [4, 5], and is further extended based on a review of recent works in which prediction models were systematically tagged and grouped. Each category of methods is discussed in a dedicated subsection, highlighting their main characteristics and limitations. For a method-by-method analysis of limitations, we refer the reader to the review [5].
@@ 4 Incorporating Driving Knowledge
Formalizing and incorporating driving knowledge into trajectory forecasting frameworks [483, 484, 485, 486] tends to enhance performance and improve the explainability of prediction models. Human drivers accumulate such knowledge through learning and experience. Similarly, forecasting models can benefit from a combination of injected driving knowledge and patterns learned from data. The intuition is that integrating external constraints narrows the space of plausible future trajectories. For example, a vehicle’s possible paths are shaped by road topology, lane directions, and adherence to traffic rules, e.g., traffic lights and right-of-way priorities at intersections and roundabouts. Vehicle motion is also constrained by physical factors such as turning radius and braking distance. Pedestrian motion, in turn, is governed by traffic rules like crossing at designated crosswalks or waiting for traffic lights at intersections. These behaviors can be formalized as structured knowledge to support trajectory forecasting. Beyond such conventional rules, less formal but still possible sources of driving knowledge include behavioral priors and region-specific driving styles. For instance, drivers may take turns from adjacent lanes to avoid congestion, or pedestrians may rush to cross even after the light has turned red, attempting to beat the incoming traffic flow due to long waiting times
Capturing these variations can enhance the performance of prediction models and improve generalizability in case the formalized knowledge is modular and can be adapted to the region of deployment. This section explores different types of driving knowledge and the ways in which this knowledge is incorporated. For a more detailed discussion, we refer the reader to the study [10].
@@ 4.1 Types of Driving Knowledge
We categorize driving knowledge into three groups: (i) map-related information that encodes the static environment, such as road topology, pedestrian crossings, stop lines, etc.; (ii) agent-related information about the kinematics of vehicles, specific driving styles and pedestrians mobility constraints due to age, disability, or behavioral factors; and (iii) traffic-related knowledge, which captures information about traffic rules and interaction patterns between road users.
@@ 4.1.1 Map-related knowledge
HD maps provide information about road geometry, traffic lights, stop lines, pedestrian crosswalks, and traffic signs. Section 2.2 discusses rasterizing map data into grid images and encoding them using CNNs. While effective for extracting spatial features, this method often results in information loss, as the structured and geometric details of the map are rasterized, reducing precision. Recent advancements use GNNs to encode HD maps, preserving structural and semantic details while remaining scalable.
One spotlight example is VectorNet, proposed by Gao et al. [73], which represents HD map elements and agent trajectories in a vectorized format. Each map element is broken into polylines composed of individual vectors, forming local graphs where nodes represent vectors and edges capture spatial or sequential relationships. These local graphs are aggregated into a global graph to encode lane-to-lane or agent-to-lane interactions using a hierarchical GNN. This approach enables interaction-aware feature encoding for downstream tasks. Next, Liang et al. [74] introduced LaneGCN, which leverages HD maps by constructing a lane graph where each lane segment is a node, and edges denote connections like predecessors, successors, and neighboring lanes. The model dynamically encodes the scene by capturing the topology and dependencies of the lane graph. By integrating actor-map interactions through a fusion network, it combines real-time traffic information with map-based context for prediction. Deo et al. [487] proposed a complementary approach, which focuses on conditioning trajectory predictions on specific paths within the lane graph. The proposed method uses discrete policy rollouts to select plausible routes, reducing computational complexity by targeting relevant map regions and decoupling lateral variability (e.g., route choices) from longitudinal variability (e.g., speed, acceleration).
@@ 4.1.2 Agent-related knowledge
Accounting for a vehicle’s kinematics is a common feature in prediction models [231, 488], as it determines turning radius, braking distance, and acceleration capacity —– all factors which influence the vehicle’s future motion. In terms of driving style, analyzing a target vehicle’s historical trajectories allows the model to associate its behavior with known driving patterns. This behavioral prior can then be used to improve prediction, e.g., an aggressive driver may be more likely to select risky maneuvers such as overtaking with insufficient gaps or forcing trailing vehicles to brake abruptly.
@@ 4.1.3 Traffic-related knowledge
Traffic-related knowledge includes additional driving information beyond agent- and map-related categories. For instance, incorporating explicit traffic rules, such as speed limits and right-of-way regulations into prediction models can help generate trajectories that better comply with expected behaviors in regulated environments. Kumar et al. [489] discuss the impact of such rules on intersection interactions, where vehicles may be required to stop, yield, or proceed depending on traffic signals or signs. In roundabouts, Hu et al. [490] highlight how right-of-way rules constrain permissible behaviors and support the generation of conflict-free trajectories for all participants. Ju et al. [491] observe that vehicle acceleration patterns often reflect compliance with traffic regulations, such as slowing near crosswalks or maintaining appropriate speeds in restricted zones.
Another source is interaction knowledge helping participants coexist in dynamic environments. In the study [489], pairwise interactions at intersections are categorized into discrete types such as ignoring, going, or yielding, based on the behavior of the target vehicle. Ban et al. [492] extend this approach by introducing follow and yield interactions, requiring models to explicitly predict these labels to better capture traffic dynamics. For more complex scenarios, such as roundabouts, Hu et al. [490] define interaction types using reference paths of interacting vehicles, enabling prediction of conflict-free trajectories. Ju et al. [491] link interactions to acceleration patterns, proposing a sub-network that predicts vehicle acceleration to infer interaction-driven trajectories. Additionally, social norms and informal yielding practices can also be of use, e.g., a driver signaling a pedestrian to pass. These implicit forms of interaction knowledge can improve prediction accuracy in heterogeneous dense traffic scenarios.